# Personalized Wireless Federated Learning for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing tasks. However, their deployment in wireless networks still face challenges, i.e., a lack of privacy and security protection mechanisms. Federated Learning (FL) has emerged as a promising approach to address these challenges. Yet, it suffers from issues including inefficient handling with big and heterogeneous data, resource-intensive training, and high communication overhead. To tackle these issues, we first compare different learning stages and their features of LLMs in wireless networks. Next, we introduce two personalized wireless federated fine-tuning methods with low communication overhead, i.e., (1) Personalized Federated Instruction Tuning (PFIT), which employs reinforcement learning to fine-tune local LLMs with diverse reward models to achieve personalization; (2) Personalized Federated Task Tuning (PFTT), which can leverage global adapters and local Low-Rank Adaptations (LoRA) to collaboratively fine-tune local LLMs, where the local LoRAs can be applied to achieve personalization without aggregation. Finally, we perform simulations to demonstrate the effectiveness of the proposed two methods and comprehensively discuss open issues.
@@ I Introduction
With the development of 6G communications, the application of Artificial Intelligence (AI) in the wireless networks is becoming increasingly important. One of the key features of 6G is the deep integration of AI with wireless networks, which can support more intelligent services and applications. Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) tasks by demonstrating impressive language understanding and generation capabilities, pushing the boundaries of AI research. LLMs can also provide more accurate understanding of user semantics and intentions, thereby offering personalized services to 6G users [1].
However, as LLM scales continue to expand, reaching hundreds of billions or even trillions of parameters, traditional publicly available datasets face challenges in meeting the demands for training future LLMs. In 6G networks, there could be a vast array of mobile devices, potentially accumulating significant volumes of user data. However, concerns regarding data security and information privacy may still prevent the users to share their personal data for the training of LLMs in wireless networks.
To leverage the vast amount of distributed and privatized data for training LLMs, Federated Learning (FL) has been adopted to offer a collaborative learning approach. This approach enables future LLMs to learn from a broader range of data sources while maintaining data security and privacy. However, there are major challenges for training the LLM by wireless FL.
@@ I-1 Big and Heterogeneous Data
LLMs require massive amounts of data from large and diverse data sources to train the model effectively. In wireless FL, the distributed structure of data is a critical challenge, as the data on each mobile device may be highly unbalanced, depending on their backgrounds, preferences, or behaviours. This can lead to slower convergence and poorer performance for training LLMs. Furthermore, diversity of data across mobile devices may introduce complex data-distribution in wireless networks.
@@ I-2 Resource-intensive Training
Training an LLM is a resource-intensive task that demands high computational power and large memory. In wireless FL, the computation is decentralized, where individual devices may need to have sufficient resources to participate in the training process. However, not all contributing devices (such as smartphones or pads) have the necessary computational and storage resources, which can lead to slow training speed or even fail to train the LLM successfully,
@@ I-3 High Communication Overhead
Wireless FL requires frequent communication between a central server and devices to update an LLM, which can incur high bandwidth and latency costs. For LLMs that have billions or even trillions of parameters, this could result in a huge amount of data to be transferred, leading to high communication overhead. Moreover, the communication cost increases with the number of communication rounds, the number of participating devices, and the model size. As such, reducing communication overhead without sacrificing model performance is a significant challenge.
Moreover, the demand for user-centric personalized LLMs has significantly increased. These LLMs have the ability to learn individual preferences and provide tailored results. Personalized Federated Learning (PFL) is an extension of FL that recognizes potential data and resource heterogeneity among different clients and aims to learn a personalized model for each client [2]. Parameter-Efficient Fine-Tuning (PEFT) enables efficient adaptation to client-specific data and tasks by adjusting a minimal number of parameters for the pre-trained LLM, thereby reducing computational resource and communication overhead. Hence, PFL combined with PEFT can overcome the aforementioned challenges associated with training LLMs by FL.
Unlike FL for generative models such as generative adversarial networks and diffusion models, LLMs require consideration of alignment with human values and interaction with external knowledge during learning. Hence, in the paper, we first introduce and compare different learning stages and their features of LLMs in wireless networks. We then summarize potential solutions of FL for various fine-tuning methods. Finally, we highlight advantages of PFL in fine-tuning LLM and propose two approaches for personalized wireless federated fine-tuning with low communication overhead as follows:
  • Personalized Federated Instruction Tuning (PFIT): We focus on Reinforcement Learning from Human Feedback (RLHF) and design two reward models to represent the helpfulness and safety by human feedback. By linearly combining the two reward models across different clients, personalized local models are obtained. Sparse self-attention is also employed to reduce communication overhead and accelerate the training speed of the federated instruction tuning.
  • Personalized Federated Task Tuning (PFTT): We combine two PEFT approaches, namely adapter and Low-Rank Adaptation (LoRA), to reduce communication overhead and accelerate the federated task fine-tuning. The adapter parameters of clients are sent to the server for global aggregation, while LoRA’s parameters are retained at the clients to maintain the personalization of local models.
The remainder of this paper can be organized as follows. Section II describes the harmonizing of LLM and FL in wireless networks. Section III provides some potential solutions of federated fine-tuning for LLMs. Section IV details the proposed personalized wireless federated fine-tuning schemes. Section V shows the simulation results. Section VI presents open issues and Section VII concludes this paper.
@@ II Harmonizing LLM and FL in Wireless Networks
Unlike traditional deep learning, LLMs have a three-stage learning process: pre-training, fine-tuning and retrieval-augmented generation (RAG). As shown in Fig. 1, the pre-training stage provides a foundation of general language understanding, while the fine-tuning stage adapts this understanding to specific tasks or goals, and the RAG stage enhances this understanding and improves the accuracy of answers by retrieving information from external data sources. All learning stages of LLMs are also summarized and compared in Table I.
Figure 1: Illustration of the three-stage learning of LLMs in wireless networks.
@@ II-A Pre-training Stage of LLM
  • High Resource Requirements: Pre-training of LLMs involves adjusting all their parameters, which requires significant computational resources. In FL, the model is trained across many devices or servers, with each device obtaining updates for the model based on its computation and storage resource [4]. Additionally, this distributed computation increases the communication overhead, as the update from each device needs to be aggregated to obtain the global model. Centralized training, on the other hand, can leverage powerful servers and optimized infrastructure to train LLMs more efficiently.
@@ II-B Fine-tuning Stage of LLM
Once the LLM completes pre-training on the server side, it undergoes further fine-tuning on a smaller, more specific dataset. This dataset typically consists of local private data. The clients download the pre-trained parameters of the LLM from the server and perform fine-tuning on an extremely small subset of parameters to enhance the performance of the LLM. To ensure security and privacy of local data, the fine-tuning process is conducted locally on the client side, and the updated subset of parameters are transmitted back to the server through wireless networks. This local dataset is often labelled, meaning that it comes with correct answers that the LLM should learn to predict. Fine-tuning requires less computational resources and smaller amount of data compared to pre-training. Fine-tuning of LLMs can be categorized into the following types:
@@ II-B1 Instruction Tuning
Instruction tuning is a strategy for fine tuning LLMs on a combination of task instructions so that the LLM can generate the correct output based on the instructions. Instruction tuning uses natural language instructions as inputs to query the LLM and guide its output. The instructions consist of sequences that pair instructions with corresponding examples, which can provide explicit and precise guidance for the LLM to generate texts that are consistent with the user’s intent and the data source. The goal of instruction learning is to improve the comprehension and generalization of LLMs on unseen tasks, as well as their helpfulness and safety.
II-B2 Task Tuning
Once the LLM has been pre-trained on the unlabelled data, it is further fine-tuned on specific downstream tasks using labelled data. Task tuning involves training the pre-trained model on a smaller task-specific dataset that is annotated with labels or target outputs. During the task tuning, the pre-trained model is adapted to the specific task by updating its parameters based on the labelled task data. The objective is to fine-tune the model’s representations and weights to better align with the target task’s objectives and improve its performance. The task tuning process helps the model generalize its pre-learned knowledge to the specific task, making it more task-specific and accurate.
Pre-training requires massive resources and utilizes publicly available data, and therefore it may not be suitable for FL. However, FL could still be valuable in the fine-tuning stage where privacy is a concern or when the tuning data is inherently decentralized. The reasons are listed as follows:
  • Low Resource Requirements: Fine-tuning is relatively efficient as it requires less data and computational resources for adjusting a small subset of parameters of LLMs compared to the pre-training, making it feasible to train LLMs even on resource-constrained devices.
@@ II-B2 Task Tuning
II-C Retrieval-Augmented Generation Stage of LLM
RAG is an approach that combines LLMs with information retrieval techniques to enhance performance of LLMs[5].
Due to the enormous training cost of LLMs, the learned knowledge has a temporal lag. For instance, the training data for GPT-3.5 is current up until January 2022, implying that it lacks knowledge of any events transpiring after January 2022. During the generation process, instead of relying solely on the pre-trained or fine-tuned model’s knowledge, retrieval mechanisms are employed to retrieve the latest relevant information from external sources such as Internet or local knowledge bases. These knowledge bases are often deployed at the edge, and clients retrieve the latest local information by querying knowledge bases. The retrieved local information, along with the client’s request, is then sent to the server-side LLM by wireless networks. Moreover, LLMs can leverage the Internet to retrieve the latest relevant public information. Then, this retrieved information is incorporated into the generated output of the LLM, ensuring that the generated content is contextually relevant and factually accurate.
However, RAG may also not be suitable for FL due to the following reasons:
  • Additional Data Exposure: Sharing sensitive retrieval queries or accessing external resources across multiple clients may compromise the privacy and security of the distributed data.
  • No Weight Update: RAG enhances the performance of LLMs by incorporating the retrieved data into the prompt for in-context learning. It does not require updating parameters of LLMs, thus eliminating the need for local gradient descent in FL.
LLMs not only need to align with human values during the learning process, but also interact with external knowledge, which represents a more complex learning paradigm. Due to the massive amount of insensitive data and extensive model parameters involved in the pre-training stage, it is more suitable for centralized cloud-based learning. In contrast, during the RAG stage, only local data embedding and querying are required, making it more suitable for local execution. Therefore, we primarily focus on designing PFLs for the fine-tuning stage.
TABLE I: Comparison of the three-stages learning of LLMs.
@@ II-C Retrieval-Augmented Generation Stage of LLM
Instruction tuning Generalization for many tasks Instruction data Supervised learning or Reinforcement learning Partial parameters (5%-10%) May require user instruction Medium computational and storage resources
Task tuning Personalization for specific tasks Task-specific data Supervised learning Few parameters (1%-2%) May require task data Low computational and storage resources
Retrieval-augmented generation Enhance output with relevant information Latest external data Retrieval No parameters Require up-to-date public or user data Low computational and high storage resources
@@ III Potential solutions of wireless federated fine-tuning for LLMs
In this section, we present key techniques and potential solutions of federated fine-tuning for LLMs.
@@ III-A Federated Instruction Tuning
Federated instruction tuning involves FL for instruction tuning of LLMs [4], which necessitates each client to possess sufficient computational resources to fine-tune a larger number of parameters compared with task fine-tuning methods. Additionally, it requires transmitting more model parameters over the network, which can be bandwidth-intensive and time-consuming. Furthermore, it is challenging to define a clear, algorithmic loss to measure the quality of the LLM’s output for local instructions. The following techniques hold the potential to address these challenges.
III-A1 Sparse Attention
In traditional attention mechanisms, every token (word) attends to every other token, resulting in a quadratic complexity with respect to the input sequence length. However, in scenarios where the input sequence is long, this approach becomes computationally expensive and memory-intensive. Sparse attention addresses this issue by introducing sparsity patterns, allowing tokens to attend only to a subset of other tokens rather than attending to all tokens [6]. This reduces the computational and memory requirements, and communication overhead while preserving the LLM’s ability to capture relevant dependencies between tokens in FL. The sparsity pattern can be learned during the instruction tuning, enabling the LLM to dynamically determine the subset of attention parameters based on the wireless environment.
III-A2 RLHF
RLHF is a Reinforcement Learning (RL)-based fine-tuning technique that improves the performance of LLMs by incorporating human-generated feedback [7] for difficult-to-assess objectives. This method involves a three-step process. Initially, an LLM is fine-tuned using supervised instruction learning, drawing from human-generated instructions or demonstrations. Next, a question and several model outputs are sampled. Human evaluators then rank these outputs from best to worst, and the ranked data is used to train a new reward model. Finally, the reward model calculates the reward for the LLM’s output, and the estimated reward is utilized to update the policy through RL. In FL, we can only adjust the parameters that are used to compute the generation probability distribution. These parameters control the probabilities of selecting different words or tokens during the generation process, which can influence the output of the LLM to better align with the expectations conveyed by human feedback (e.g., to fine-tune only the last few layers of LLMs).
III-B Federated Task Tuning
@@
Federated task tuning employs FL for fine-tuning downstream tasks of LLMs [8], which is relatively efficient as it requires less data and computational resources of cilents than that of the instruction tuning. It allows for task-specific adaptation but also risks overfitting or forgetting the original knowledge of local LLMs. PEFT is a set of techniques designed to adapt LLMs to specific downstream tasks with minimal changes to the original parameters. The following PEFT methods can help mitigate overfitting and catastrophic forgetting in LLMs.
III-B1 Adapter
Adapter tuning is introduced to add small, task-specific adapter modules to a pre-trained LLM [9]. These adapter modules are inserted between the layers of the pre-trained LLM. They typically consist of a bottleneck structure: a down-projection, a non-linearity, and an up-projection. During fine-tuning, only the parameters of these adapter modules are updated, while the parameters of the pre-trained LLM are kept fixed. In FL, this approach significantly reduces the number of parameters that need to be updated in each client, while still allowing the LLM to adapt to the specific task with fewer resources. Moreover, when adaptation to wireless channel quality, we can define the dimensions of adapters adaptively, thereby dynamically adjusting the communication overhead for transmission in wireless channels.
III-B2 LoRA
@@
LoRA adapts a pre-trained LLM to a new task by applying low-rank matrix decomposition to the model parameters [8]. LoRA decomposes the original weight matrices into the product of two smaller matrices, and only updates the smaller matrices during fine-tuning. In FL, LoRA can reduce the number of parameters that need to be updated, thereby decreasing the communication and computational cost on the client. By preserving historical low-rank matrices, LoRA can prevent catastrophic forgetting of local models, while preserving its generation and generalization abilities.
Federated instruction tuning deepens the understanding of queries by learning from human instructions, while federated task tuning enhances the performance of downstream tasks by learning from different task data. Both approaches strengthen the distributed learning performance of LLMs from different perspectives.
IV Personalized federated fine-tuning

### v2

@@ Abstract
Large language models (LLMs) have driven profound transformations in wireless networks. However, within wireless environments, the training of LLMs faces significant challenges related to security and privacy. Federated Learning (FL), with its decentralized architecture, offers enhanced data privacy protection. Nevertheless, when integrated with LLMs, FL still struggles with several critical limitations, including large-scale and heterogeneous data, resource-intensive training, and substantial communication overhead. To address these challenges, this paper first presents a systematic analysis of the distinct training stages of LLMs in wireless networks, including pre-training, instruction tuning, and alignment tuning. Building upon this foundation, we propose a Personalized Wireless Federated Fine-tuning (PWFF) framework. Initially, we utilize the adapter and Low-Rank Adaptation (LoRA) techniques to decrease energy consumption, while employing global partial aggregation to reduce communication delay. Subsequently, we develop two reward models and design a personalized loss function to fulfill the goal of personalized learning. Furthermore, we implement a local multi-objective alignment to ensure the stability and effectiveness of the FL process. Finally, we conduct a series of simulations to validate the performance of the proposed PWFF method and provide an in-depth discussion of the open issues.
@@ I Introduction
With the exploration of future 6G communication paradigms, the application of Artificial Intelligence (AI) in wireless networks is becoming increasingly important. One of the key features of 6G is the deep integration of AI with wireless networks, which can support more intelligent services and applications. Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) tasks by demonstrating impressive language understanding and generation capabilities, pushing the boundaries of AI research. LLMs can also provide a more accurate understanding of user semantics and intentions, thereby offering personalized services to 6G users [1].
However, as LLM scales continue to expand, reaching hundreds of billions or even trillions of parameters, traditional publicly available datasets face challenges in meeting the demands for training future LLMs. In 6G networks, there could be a vast array of mobile devices, potentially accumulating significant volumes of user data. However, concerns regarding data security and information privacy may still prevent users from sharing their personal data for the training of LLMs in wireless networks.
To leverage the vast amount of distributed and privatized data for training LLMs, Federated Learning (FL) has been adopted to offer a collaborative learning approach. This approach enables future LLMs to learn from a broader range of data sources while maintaining data security and privacy. However, there are major challenges to training the LLM by wireless FL.
@@ I-1 Big and Heterogeneous Data
LLMs require massive amounts of data from large and diverse data sources to train the model effectively. In wireless FL, the distributed structure of data is a critical challenge, as the data on each mobile device may be highly unbalanced, depending on their backgrounds, preferences, or behaviors. This can lead to slower convergence and poorer performance for training LLMs. Furthermore, the diversity of data across mobile devices may introduce complex data distribution in wireless networks.
@@ I-2 Resource-intensive Training
Training an LLM is a resource-intensive task that demands high computational power and large memory. In wireless FL, the computation is decentralized, where individual devices may need to have sufficient resources to participate in the training process. However, not all contributing devices (such as smartphones or pads) have the necessary computational and storage resources, which can lead to slow training speed or even to suboptimal performance.
@@ I-3 High Communication Overhead
Wireless FL requires frequent communication between a central server and devices to update an LLM, which can incur high bandwidth and latency costs. For LLMs that have billions or even trillions of parameters, this could result in a huge amount of data being transferred, leading to high communication overhead [2]. Moreover, the communication cost increases with the number of communication rounds, and the number of participating devices. As such, reducing communication overhead without sacrificing model performance is a significant challenge.
Moreover, the demand for user-centric personalized LLMs has significantly increased [3]. These LLMs have the ability to learn individual preferences and provide tailored results. Personalized Federated Learning (PFL) is an extension of FL that recognizes potential data and resource heterogeneity among different clients and aims to learn a personalized model for each client [4]. Parameter Efficient Fine-Tuning (PEFT) enables efficient adaptation to client-specific data and tasks by adjusting a minimal number of parameters for the pre-trained LLM, thereby reducing computational resources and communication overhead [2]. Hence, PFL combined with PEFT can overcome the aforementioned challenges associated with training LLMs by FL.
In this paper, we investigate personalized wireless FL for LLMs. First, we introduce the two learning phases in LLMs: pre-training and fine-tuning, analyzing the advantages and disadvantages of conducting FL in wireless networks during each phase. Next, we discuss several challenges associated with the federated fine-tuning phase and potential solutions. We then analyze the benefits of PFL and propose a Personalized Wireless Federated Fine-tuning (PWFF) framework for LLMs. In this framework, we utilize the adapter and Low-Rank Adaptation (LoRA) techniques to decrease energy consumption, while employing global partial aggregation to reduce communication delay. Subsequently, we develop two reward models and design a personalized loss function to fulfill the goal of personalized learning. Finally, we introduce a local multi-objective alignment to ensure the stability and effectiveness of the learning process for FL. Experimental results demonstrate the advantages of our proposed method in terms of communication latency and energy consumption.
The remainder of this paper can be organized as follows. Section II describes the harmonizing of LLM and FL in wireless networks. Section III provides some potential solutions for federated fine-tuning for LLMs. Section IV details the proposed PWFF scheme. Section V shows the simulation results. Section VI presents open issues, and Section VII concludes this paper.
@@ II Harmonizing LLM and FL in Wireless Networks
Unlike traditional deep learning, LLMs have a two-stage learning process: pre-training and fine-tuning, as shown in Fig. 1. The pre-training stage provides a foundation of general language understanding, while the fine-tuning stage adapts this understanding to specific tasks or goals. All learning stages of LLMs are summarized and compared in Table I.
Figure 1: Illustration of the two-stage learning process of LLMs in wireless networks.
@@ II-A Pre-training Stage of LLM
  • High Resource Requirements: Pre-training of LLMs involves adjusting all their parameters, which requires significant computational resources. In FL, the model is trained across many devices or servers, with each device obtaining updates for the model based on its computation and storage resources. Additionally, this distributed computation increases the communication overhead, as the update from each device needs to be aggregated to obtain the global model. Centralized training, on the other hand, can leverage powerful servers and optimized infrastructure to train LLMs more efficiently.
@@ II-B Fine-tuning Stage of LLM
Once the LLM completes pre-training on the server side, it undergoes further fine-tuning on a smaller, more specific dataset. This dataset typically consists of local private data. The clients download the pre-trained parameters of the LLM from the server and perform fine-tuning on an extremely small subset of parameters to enhance the performance of the LLM. To ensure the security and privacy of local data, the fine-tuning process is conducted locally on the client side, and the updated subset of parameters is transmitted back to the server through wireless networks. This local dataset is often labelled, meaning that it comes with correct answers that the LLM should learn to predict. Fine-tuning requires fewer computational resources and a smaller amount of data compared to pre-training. Fine-tuning of LLMs can be categorized into the following types:
@@ II-B1 Instruction Tuning
Instruction tuning is a strategy for fine-tuning LLMs on an instruction dataset so that the LLM is capable of better understanding instructions and generating accurate outputs based on the corresponding instructions. This approach aims to enhance the model’s ability to follow user instructions across a wide range of tasks by aligning its responses with the intended outcome of those instructions. The process of instruction tuning uses supervised learning on the instruction dataset. By iteratively updating the LLM’s parameters through back-propagation, the LLM learns to generalize across various instruction types.
II-B2 Alignment Tuning
Alignment tuning is the process of optimizing the LLM to ensure that its outputs align with human values, preferences, and ethical norms. This approach seeks to make the LLM’s responses more safe, reliable, and beneficial by reducing harmful or undesirable outputs. This process strengthens positive behaviors through reward mechanisms while discouraging negative ones, ultimately aligning the LLM with human intentions, and avoiding harmful, inaccurate, or unethical content.
FL could be valuable in the fine-tuning stage where privacy is a concern or when the tuning data is inherently decentralized. The reasons are listed as follows:
  • Low Resource Requirements: Fine-tuning is relatively efficient as it requires less data and computational resources for adjusting a small subset of parameters of LLMs compared to pre-training, making it feasible to train LLMs even on resource-constrained devices.
@@ II-B2 Alignment Tuning
LLMs not only need to acquire extensive knowledge but also require alignment with human values during the learning process, which represents a more complex learning paradigm. Due to the massive amount of insensitive data and extensive model parameters involved in the pre-training stage, it is more suitable for centralized cloud-based learning. Therefore, we primarily focus on designing PFLs for the fine-tuning stage.
TABLE I: Comparison of the two-stage learning process of LLMs.
@@
Instruction tuning Understanding user instructions Instruction data Supervised learning Few parameters May require user instructions Low computational and storage resources
Alignment tuning Aligning with user preferences Preference-specific data Reinforcement learning Few parameters May require user preferences Low computational and storage resources
@@ III-A Federated Instruction Tuning
Federated instruction tuning is a federated supervised learning for LLMs. It allows for instruction-specific adaptation but also risks overfitting or forgetting the original knowledge of local LLMs [6]. PEFT is a set of techniques designed to adapt LLMs to specific instructions with minimal changes to the original parameters [2]. The following PEFT methods can help mitigate overfitting and catastrophic forgetting in LLMs.
III-A1 Adapter
Adapter tuning achieves task-specific parameter updates by inserting lightweight adapter modules between the layers of a pre-trained LLM [7]. These adapters typically adopt a bottleneck architecture, consisting of a down-projection layer, a non-linear activation function, and an up-projection layer. During fine-tuning, only the parameters within the adapter modules are updated, while the core weights of the pre-trained LLM remain fixed, significantly reducing the number of trainable parameters. In FL, this approach lowers the computational and storage overhead of local training, making it particularly suitable for personalized fine-tuning in resource-constrained environments.
III-A2 LoRA
@@
LoRA adapts a pre-trained LLM to a new task by applying low-rank matrix decomposition to the model parameters [8]. LoRA decomposes the original weight matrices into the product of two smaller matrices, and only updates the smaller matrices during fine-tuning. In FL, LoRA can reduce the number of parameters that need to be updated, thereby decreasing the communication and computational cost of the client. By preserving historical low-rank matrices, LoRA can prevent catastrophic forgetting of local models, while preserving its generation and generalization abilities.
III-B Federated Alignment Tuning
Federated alignment tuning is a federated Reinforcement Learning (RL) for LLMs[8]. The LLM is fine-tuned by using human feedback to generate reward signals that guide the model’s behavior. However, it is challenging to define a clear, algorithmic reward to measure the quality of the LLM’s output for local user preferences. The following techniques hold the potential to address these challenges.
III-B1 Reinforcement Learning from Human Feedback (RLHF)
RLHF is a fine-tuning method designed to enhance the alignment of LLMs with human preferences and task-specific expectations [9]. The LLM generates multiple candidate responses based on given prompts, which are then ranked by human annotators. These ranked responses are used to train a reward model that quantifies the degree of alignment between the model outputs and human preferences. In the RL stage, the reward model guides the optimization of the LLM’s policy. In FL, RLHF can be adapted by updating only a small subset of critical parameters (e.g., LoRA), thereby enabling efficient alignment with human feedback across distributed clients.
III-B2 Multi-objective preference alignment
Although LLMs can be aligned with collective average preferences through RLHF, such models often fail to satisfy diverse individual preferences [10]. Each person places different emphasis on various alignment dimensions (e.g., usefulness, harmlessness, safety). To achieve multi-objective alignment, a weighted combination of multiple reward functions can be used to represent individual preferences, thereby enabling multi-objective RL. However, in federated learning, the training process of multi-objective RL is complex and unstable, requiring the use of specialized robust learning methods to enhance the performance of multi-objective federated RL.
IV Personalized Wireless Federated Fine-tuning